A Restaurant manager decides to lower the price of a featured sandwich from Rs. 500 to Rs. 400, and she finds that sales during the week increase from 440 to 680 sandwiches. Is demand elastic or Inelastic?


1
Expert's answer
2021-05-03T10:58:09-0400

Solution:

Price elasticity of demand (PED) ="\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"


% change in qty demanded ="\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \times 100"


= "\frac{680 -440}{(680+440)/2 } \times 100 = \frac{240}{560} \times 100 = 42.86\%"


% change in price = "\frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 } \times 100"


= "\frac{400 -500}{(400+500)/2 } \times 100 = \frac{-100}{450} \times 100 = -22.22\%"


Price elasticity of demand (PED) = "\frac{42.86\%}{-22.22\%} = -1.93"

PED = 1.93

The demand is elastic. This is because PED is greater than one, meaning that consumers are very sensitive to sandwich price changes.